OpenStudy (will.h):

@Pooja195 Help ASAP Medal and everything :)

OpenStudy (will.h):

Hi i attached the whole file so you can take a look at what i've done in the previous questions because all questions are connected somehow i need help with the last question

OpenStudy (will.h):

@AloneS

OpenStudy (will.h):

Hi @phi can you help pls

OpenStudy (phi):

I'm not sure about this. but I would guess you start with 1+ 0.03/12 = 1.0025 as the monthly interest rate then they say find \[ 1.0025^\frac{1}{4}= 1.000624415 \] using a calculator the adjusted rate is the "decimal part" .000624415 * 48 (instead of 12) = 0.02997 or 2.997%

OpenStudy (will.h):

@phi doesn't make any sense because the function is already decay! so here's what i did after i reached this point 10,000 = 39,145(1.0025^1/4)^48t i simplified what's inside the parentheses and then 0.224 = 1 + r -0.776 = r i don't know am not sure at all w this one!

OpenStudy (phi):

What value did you use for "t" ?

OpenStudy (will.h):

i didn't solve for t kept it the way it is. i just simplified what's inside the parentheses and 1/4 is the exponent means that for 1 month there are 4 weeks so then i found the r.

OpenStudy (phi):

I tried to answer this same question a few years ago http://openstudy.com/study#/updates/53694306e4b0ff6f8a496ff8 Here is the last post in that question Hey @phi I called my math teacher and he said that all it was asking to do was replacing "n" with the numbers of weeks instead of months. So instead of n=12 it would be n=52. Then solve like we did in my last question we did together.

OpenStudy (will.h):

my teacher will be on in 1 hr i shall ask him and tell you if you were correct. Thanks phi

OpenStudy (phi):

ok good luck

OpenStudy (mathmate):

It seems to me that something (perhaps the first parts of the question) is missing. For example, we do not know what the monthly payment is. That is why the answer is -45 months, meaning that if you borrowed 10000 45.5 months ago, you would owe them $39145 today. This is not normal, according to the supplied formula, which is just plain compound interest, without relation to monthly payments. Also, if you decide to monitor payments 4 times a month, the interest is still compounded monthly, not 48 times a year (according to the supplied formula). I think the person who designed the question either does not know math at all, or does not know (or care about) how things work in reality. Sigh!

OpenStudy (will.h):

@mathmate what you refereed to is smart tbh in part 2 the number of months turned to negative! -45! how does that even work even though if in decrease the principal as they noted! time can't be negative they said that in their lessons a million time lol